1. If the standard deviations for return on stock A and stock B are 28% and 26%

respectively and the betas for two stock are 0. 8 and 0.9, which statement as following is

correct?

A. Stock A has a higher systematic risk than Stock B

B. Stock A has a lower total risk than Stock B

C. Stock A has a higher total risk than Stock B

D. All of the above statements are false

2. You purchased XYZ Inc common stock one year ago for $50 per share. You received a

dividend of $2 per share today and decide to take your profits by selling at $54.50 per

share. What are your holding period return and capital gain yield respectively?

A)13.0%and 9%

B)13.0%and 3.5%

C)9%and 3.5%

D)9%and 13%
Answer #1

1.

As beta of stock A is lower this means it has lower systematic risk,

But Standard Deviation of Stock A is higher which means it has higher total risk.

Option C is correct.

2.

HPY = (2 + (54.50 - 50)/50 = 13%

Capital Gain Yield = 4.50/50 = 9%

Option A is correct.